An economy has no government sector and no international trade. One year, desired investment spending equals $481 billion, consumption spending in this economy equals $623 billion, and total output equals $1,401 billion.

4.1 Is this economy in equilibrium? Explain why or why not.

4.2 What is the unplanned increase in inventories in this economy?

4.3 What is total investment in this economy?
Answer #1

4.1)

Answer: not in equilibrium

(All figures are in billion dollars)

Total output = Y = 1,401

Consumption = C = 623

Desired or planned investment = Ip = 481

Government spending = G = 0

Net export = Nx = 0

This will be in equilibrium if the following equation is satisfied.

Y = C + Ip + G + Nx

1,401 = 623 + 481 + 0 + 0

1,401 = 1,104

Since left side and right side of the equation are not equal, the economy is not in equilibrium.

4.2)

Answer: 297

Since Y is greater than (C + Ip + G + Nx), there is an increase in unplanned inventory.

Increase in unplanned inventory = Y – (C + Ip + G + Nx)

                                                     = 1,401 – 1,104

                                                     = 297

4.3)

Answer: 778

Actual or total investment (I) must be equal to Ip and unplanned inventory changes.

I = Ip + Inventory changes

= 481 + 297

= 778